A civil engineer takes a public ride going to his firm. The average time for his trip is 24 minutes, with a standard deviation of 3.80 minutes. If the office offers coffe from 8:50AM to 9:00AM, and he leaves his house at 8:35AM, what is the probability (in percent) that he misses coffee?


1
Expert's answer
2021-05-18T02:19:02-0400

Required Probability

P(X>25)=1-P(X<=25)

=1-P((X-mu)/sigma<=(25-24)/3.8)

=1-(P(Z)<=0.23)

=1-0.6026

=0.3974